An ideal gas has a volume of 100cm3³ at 1.1*10^5Pa and 27°C. What is it volume at 2*10^5Pa and 60°C?
1
Expert's answer
2019-04-15T10:04:26-0400

The equation of state of ideal gas is "p V = \nu R T", where "p" is pressure, "V" is volume, "\nu" is the number of moles, "T" is the absolute temperature, and "R" is the universal gas constant. From this equation, we obtain the ratio of volumes "V_1" and "V_2" at different corresponding values of pressure and absolute temperature "p_1, T_1" and "p_2, T_2":

"\frac{V_2}{V_1} = \frac{p_1}{p_2} \frac{T_2}{T_1} \, ."


According to the conditions, we have "V_1 = 100\, \text{cm}^3", "p_1 = 1.1 \times 10^5\, \text{Pa}", "T_1 = (27 + 273.15)\, \text{K} = 300.15\, \text{K}", "p_2 = 2 \times 10^5\, \text{Pa}", and "T_2 = (60 + 273.15)\, \text{K} = 333.15\, \text{K}". Then "p_1 \/ p_2 = 0.55" and "T_2 / T_1 \approx 1.11", and we obtain

"V_2 \approx 0.55 \times 1.11 V_1 \approx 0.61 \times 100\, \text{cm}^3 = 61 \, \text{cm}^3 \, ."



Answer: 61 cm3.